OpenStudy (anonymous):

19. Write the equation of a line with slope m = 6 and including point (0, 1). A. y=6x-1/2 B. y=6x+1/2 C. y = 6x + 1 D. y = 6x - 1

OpenStudy (lxelle):

use equation y1-y2=m(x1-x2)

OpenStudy (lxelle):

y-1=6(x-0)

OpenStudy (dangerousjesse):

Y=6x+1 :)